f(x)=1/3^x raised to the x power for x values -3, -2, -1, 0, 1, 2, 3
Answer by stanbon(57967)

f(x)=(1/3)^x raised to the x power for x values -3, -2, -1, 0, 1, 2, 3
---------------
Note: The negative on the exponent is an instruction; it says "invert
the base".
---------------------------------
f(-3) = (1/3)^(-3) = 3^3 = 27
f(-2) = (1/3)^(-2) = 3^2 = 9
f(-1) = (1/3)^-1 = 3
f(0) = (1/3)^x = 1
f(1) = (1/3)^1 = 1/3
f(2) = (1/3)^2 = 1/9
f(3) = (1/3)^3 = 1/27
======================
Cheers,
Stan H.